t2113

Hobart 12m:

Disk VSN: HOB+0098 (7099.712 GB)

  • 17:30 UT - Experiment started okay (Ross)

Katherine 12m:

Disk VSN: HOB+0122 (14343.809 GB) (Part 1/2, together with r1762)

        USN-0067 (0.0 GB) (Part 2/2)
        
* 17:30 UT - Experiment started okay (Ross)
* 2325 UT - Experiment halted to swap banks to allow R1 shipment. New module is USN-0067/4000/1024 (Initial disk_pos 0.0) (JMc)
* 2337 UT - Experiment resumed, first good data. All data between 2325 and 2336 unusable. (JMc)
* 0050 UT - MET sensor disconnected - GA upgrading GNSS receiver & new interface is required. (JMc)

Yarragadee 12m:

Disk VSN: BKG-0015 (0.000 GB)

  • 17:30 UT - Experiment started okay (Ross)

~~DISCUSSION~~